OpenStudy (anonymous):

x^2+3x-9=0 completing the square. please will appreciate if working is shown.

OpenStudy (triciaal):

a = 1 b = 3 and c = -9 what is -b/2a =

OpenStudy (triciaal):

can you respond?

OpenStudy (anonymous):

sorry just got back

OpenStudy (dumbcow):

let me explain with general equation and see of that makes sense to you. starting with \[x^2 +bx+c=0\] you want the equation to end up looking like \[(x+h)^2 + k = 0\] expanding the square: \[x^2 +2hx +h^2 +k = 0\] this means \[b = 2h \rightarrow h = \frac{b}{2}\] and \[c = h^2 +k \rightarrow k = c - h^2\]

OpenStudy (dumbcow):

For your example; b= 3 c = -9 so \[h = \frac{b}{2} = \frac{3}{2}\] \[k = c - h^2 = -9 - (\frac{3}{2})^2\]